OpenStudy (anonymous):

If u = 1 + i and v = 1 – i, what is u + v? a)2 b)0 c)-2i d)-2 e)2i

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

c?

OpenStudy (michele_laino):

here are some steps: \[\Large \begin{gathered} u + v = \left( {1 + i} \right) + \left( {1 - i} \right) = 1 + i + 1 - i = \hfill \\ \hfill \\ = \left( {1 + 1} \right) + i\left( {1 - 1} \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

option C) is a wrong option

OpenStudy (anonymous):

yes i know i got it thank you :)

OpenStudy (michele_laino):

:)